Can someone please help me with this problem. Please, please.
When playing bingo, 75 balls are placed in a bin and balls are selected at random. Each ball is marked with a letter and number as indicated in the following chart.
B I N G O
1-15 16-30 31-45 46-60 61-75
For example, there are balls marked B1, B2 up to B15; I16, I17, up to I30; and so on. Assuming one bingo ball is selected at random, determine
the odds in favor of it being B9

Answer by sudhanshu_kmr(1152)


probability that ball is B9 i.e P(A)= 1/75
probability that ball is not B9 i.e P'(A)= 74/75

odds in favor of it being B9 = P(A)/P'(A) = 1/74


If any doubt, you are welcome to get free help..